A shopper buys three oranges and five lemons for $10.26, while a second shopper buys four lemons and six oranges for $11.16. What is the price of each fruit?
Answer by checkley71(8403)

3O+5L=$10.26 & 6O+4L=$11.16 THUS 3O=10.26-5L THEREFORE 6O IN THE SECOND EQUATION =2(10.26-5L) OR 20.52-10L OR 20.52-10L+4L=11.16 OR -6L=-9.36 OR L=9.36/6 OR L=$1.56 THEN 3O+5*1.56=10.26 OR 3O+7.80=10.26 OR 3O=2.46 OR O=2.46/3 OR O=$.82